OpenStudy (anonymous):

if the angles A ,B ,C of a triangle are in AP and if a,b,c denotes the lenght of the side opposite to A, B , C respectively then the value of the expression a/c *sin2C + C/a * Sin2A is ?

hartnn (hartnn):

A+B+C=180 A+C=2B B=.... ?

OpenStudy (anonymous):

B=60

OpenStudy (shubhamsrg):

how about hit and trial.. assume the angles are 60,60,60 so final ans becomes sqrt3 ! :P please confirm..

OpenStudy (anonymous):

Yes..)

OpenStudy (anonymous):

but i did nt get u..)

hartnn (hartnn):

from sine rule, c/a =... ? then c/a *(2sin A cos A )= .. ?

OpenStudy (anonymous):

Ha...Silly...i got it..:)

hartnn (hartnn):

wasn't it easy ? ;)

OpenStudy (raden):

what means AP ?

hartnn (hartnn):

Arithmetic Progression if, a,b,c are in AP, then 2b=a+c

OpenStudy (raden):

actually, if a,b,c in AP then the triangle have sides 3x, 4x, 5x. it means the right triangle. right ?

hartnn (hartnn):

how did u come up with sides 3x, 4x, 5x ? and angels are in AP, not sides....

OpenStudy (raden):

OPSS... i look ur coment before hartnn :)